Fermi National Laboratory
 

Accelerator Update for July 7, 2006 - July 10, 2006

 

Friday July 7

The day shift began with Operations monitoring store 4815 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with NuMI off due to a horn cooling problem, and with MiniBooNE taking beam.

Operations established beam to Switchyard (SY) and to the Meson MTest experiment at 9:34 AM.

Operations began preparing the Recycler at 12:01 PM to stash antiprotons. The stash ended and Pbar resumed stacking at 1:04 PM.

MTest resumed beam at 1:39 PM.

MTest experimenters reported that they were done with the beam. Operations disabled their beam at 9:02 PM.

The Recycler began preparing to stash at 9:10 PM. The stash was delayed while operators tried to resolve data communication problems with the Accumulator gap monitor (GPIB errors). The stash ended and Pbar resumed stacking at 10:35 PM.

Saturday July 8

The midnight shift began with Operations monitoring store 4815 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

Operations terminated store 4815 at 2:02 AM and then put the TeV into shot setup fourteen minutes later.

Operations established store 4816 at 4:30 AM with an initial luminosity of 140.17E30.

The Recycler began preparing to stash at 4:59 AM. The stash ended and Pbar resumed stacking at 5:20 AM.

The Recycler began preparing to stash at 8:36 AM. The stash ended at 9 AM. Operators reported that there were more GPIB errors. Operations contacted experts.

Operations established beam to MTest at 9:06 AM.

Around 12:09 PM, a TeV RF station (TRF1) tripped off, which caused a loss of luminosity. Operations contacted an expert.

A Booster expert began a study at 12:26 PM.

The Recycler began preparing to stash at 3:33 PM.

At 4:25 PM, a Booster kicker (MKS05) tripped off. An operator turned it on locally.

The stash ended and Pbar resumed stacking at 4:46 PM.

MTest finished using the beam at 9:03 PM.

The Recycler began preparing to stash at 9:34 PM. The stash ended and Pbar resumed stacking at 9:59 PM.

At 11:30 PM, Operations reported that the Booster 400MeV area chopper was missing pulses.

Sunday July 9

The midnight shift began with Operations monitoring store 4816 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 12:02 AM, the 400 MeV area chopper was missing pulses again. Operations contacted an expert and then turned off the chopper.

At 12:48 AM, Operations reported that a number of TeV vacuum pumps tripped off and two cryo system valves had closed. The cryo system, though oscillating, was still operating.

At 1:05 AM, Operations reported that the cryo system was no longer operating normally. The vacuum pumps would not reset. Cryo system technicians at CHL began trying to stabilize the system. Operations advised CDF and D0 to turn off their detectors.

The Recycler began preparing to stash at 2:17 AM. The stash ended and Pbar resumed stacking at 2:38 AM.

Operations reported at 2:23 AM that the cryo system, though still oscillating, was stable enough for the experiments to turn back on.

At 4:26 AM, Operations told CDF and D0 that the cryo system was going unstable and that they should put their detectors in standby.

At 4:33 AM, Operations terminated store 4816 before there was a quench. Operations called in vacuum experts and prepared the TeV for an access.

Operations prepared D0 for an access at 5:39 AM.

The Recycler began preparing to stash at 6:57 AM. The stash ended and Pbar resumed stacking at 7:17 AM.

TeV vacuum experts reported at 9:19 AM that they had replaced vacuum pumps at TeV sectors E13, D17 and 47, and performed maintenance on a pump at D37. They checked for vacuum leaks, but found none. Operations began the TeV checkout.

Operations reported at 10:53 AM that the vacuum and cryo systems at sector E1 were recovering.

The Recycler began preparing to stash at 10:44 AM.

At 11:31 AM, a Booster kicker (MKS05) tripped off. Operators had the kicker power supply back on seventeen minutes later.

The stash ended and Pbar resumed stacking around 11:42 AM. Operations had prepared MTest for beam by 12:02 PM.

At 12:30 PM, Operations reported that the vacuum recovery had slowed. Operation contacted an expert who said that techs would come back in and install more vacuum pumps.

Operations prepared the TeV at 2:10 PM for another access and vacuum techs installed two more vacuum pumps. They came out of access at 4:42 PM. The vacuum improved and operators began the TeV turn on.

MTest resumed taking beam at 4:56 PM.

Operations put the TeV through a dry squeeze at 5:48 PM.

Operations put the TeV into shot setup at 6:14 PM.

At 6:45 PM, Operations reported that the TeV shot was on hold due to a problem with a Front End. An operator rebooted the front end per expert instruction. The shot setup was resumed at 7:01 PM.

Operations established store 4818 at 9:07 PM with an initial luminosity of 152.61E30.

The Recycler prepared to stash at 9:43 PM. The stash ended at 10:10 PM. Pbar resumed stacking.

Monday July 10

The midnight shift began with Operations monitoring store 4818 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler prepared to stash at 2:30 AM. The stash ended and Pbar resumed stacking by 2:50 AM.

At 5:04 AM, Operations reported a Booster kicker Klixon temperature sensor fault for MKS09. Operations called in an expert.

At 7:16 AM, Operations reported that the Pbar stack was unstacking. Operations investigated and discovered a bad PLC for the transverse core cooling in the Accumulator.

At 7:33 AM, Operations immediately began to stash antiprotons. It was an inefficient transfer, but better than losing all the antiprotons.

Operations began preparing Pbar at 7:58 AM for an access.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/10/2006    email Fermilab